# Model Predictive Inferential Control of Neural State-Space Models for Autonomous Vehicle Motion Planning

### v4

@@ Abstract
Index Terms:
Model predictive inferential control, model predictive control, motion planning, implicit importance sampling, neural-state-space model, particle filtering, particle smoothing.
@@ II-B Optimal Control via Estimation
The connections between control and estimation have been a fundamental research topic. The seminal work [35] by Rudolf Kalman in 1960 elucidates the duality between the linear quadratic regulator and Kalman smoother for linear systems. Even though such exact duality was long considered hardly generalizable, some illuminating studies manage to formulate nonlinear stochastic optimal control and estimation problems dual to each other [36, 37, 38]. Optimal control computation often requires great amounts of time and memory, but the control-estimation synergy makes it possible to cast control design as explainable estimation problems which are more tractable to solve. The work in [39] proposes stochastic optimal control by variational inference based on Kullback-Leibler divergence minimization, which has initiated a line of inquiry. Especially, a special case of the approach [39] is the path integral control method developed in [40]. This relevance has inspired a body of work known as model predictive path integral control (MPPI) [41, 42]. One can use numerical variational inference methods based on stochastic optimization, Markov Chain Monte Carlo sampling, importance sampling, and others, to compute optimal control for these methods. Further, stochastic optimal control is handled in [43, 44] through recursive Bayesian state estimation based on belief propagation; in [45], it is translated into a parameter inference problem by introducing a parameterized control policy and then addressed via an expectation-maximization-based inference procedure. For the studies [43, 44, 45], only approximate relationships exist between the respective control and estimation problems. However, for deterministic nonlinear systems, optimal control and MPC problems can find equivalent counterparts in recursive Bayesian state estimation problems [46], and particle filtering provides a useful means to deal with these problems due to its power in handling nonlinearity. In our preliminary work [47, 48], we develop constraint-aware particle filtering/smoothing to perform nonlinear MPC with constraints. Closely tied to optimal control, reinforcement learning has found some solutions based on Bayesian estimation [49] and variational inference [50].
@@ II-C MPC of Machine Learning Dynamic Models
The sweeping successes of machine learning have exponentially dtriven the study of learning-based MPC for dynamic systems [51, 52]. A vast literature has formed around a rich set of topics. We narrow the attention here to the specific subject of MPC of neural networks. Neural networks have a history of being used in the data-driven modeling of dynamic systems. Such models are often called neural state-space (NSS) models, and MPCs for them have garnered many applications [52]. To date, the solvers are mostly based on gradient-based optimization [53, 54, 55, 56, 57, 58]. Gradient-based search, however, will easily lead MPC computation to get stuck in local minima, as NSS models are highly nonlinear and nonconvex. The study in [59] suggests explicitly constructing neural networks that are convex with respect to the input so as to avoid nonconvex optimization, but such input-convex neural networks would have restricted representation capacity. MPC of NSS models is an essential step in model-based reinforcement learning but faces unaffordable computation when applying gradient-based solutions to models based on deep neural networks. In this context, a simple technique is to randomly generate many candidate control decision sequences and then pick the sequence that leads to the minimum cost after being applied to the NSS model [6, 60]. This technique, however, can hardly balance between accuracy and computation. The problem of MPC for NSS models hence still remains widely open to new solutions. On a related note, MPC for systems described by Gaussian processes has recently seen burgeoning studies, e.g.,  [61, 62, 5], and there arise similar challenges.
@@ III-A Neural State-Space Vehicle Modeling
We can construct $\hat{f}$ using different numerical discretization methods. A straightforward way is using the first-order Euler method:
@@ III-B Driving Requirements and Objectives
An autonomous vehicle is expected to operate responsibly within traffic, demonstrating safety, ethics, and predictability in its maneuvers. To this end, it should adhere to some driving requirements and constraints. Specifically, the vehicle (referred to as the ego vehicle or EV) must avoid collision with other traffic participants (referred to as obstacle vehicles or OVs), remain within the boundaries of the road, operate in its practical actuation limits, and follow some nominal guidance (in path, velocity, etc.) from a higher-level decision maker.
@@
$\mathcal{S}_{k}=\mathcal{S}(\bm{x}_{k},\mathcal{B}),\ \ \mathcal{O}_{k}^{i}=\mathcal{O}\left(\bm{x}_{k}^{\mathrm{OV},i},\mathcal{B}\right),\ i=1,2,\ldots,N_{O}.$
@@
Vehicle Actuation Limits: The EV’s actuation $\bm{u}_{k}$, including the acceleration and steering angle, is subject to physical limits during the maneuver. We thus constrain $\bm{u}_{k}$ as follows:
@@ III-C Motion Planning Synthesis
Having laid out the driving requirements and objectives, we are now ready to synthesize the motion planning problem. To begin with, we express the constraints in (2)-(5) compactly as
@@
$\min\quad$ $J(\bm{x}_{k:k+H},\bm{u}_{k:k+H},\Delta\bm{u}_{k:k+H}),$ (8a)
@@ IV-A MPC Through the Lens of State Estimation
[…]
@@ IV-B MPIC Based on Bayesian State Estimation
$\bar{\bm{f}}(\bar{\bm{x}}_{t})=\begin{bmatrix}\hat{f}(\bm{x}_{t},\bm{u}_{t})\\ \bm{u}_{t}\\ \bm{0}\end{bmatrix},\;\bar{\bm{h}}(\bar{\bm{x}}_{t})=\left[\begin{matrix}\bm{x}_{t}\cr\bm{u}_{t}\cr\displaystyle\sum_{j=1}^{m}\psi(g_{j}(\bm{x}_{t},\bm{u}_{t},\Delta\bm{u}_{t}))\end{matrix}\right].$
@@
We will use $p(\bar{\bm{x}}_{k:k+H}\,|\,\bar{\bm{y}}_{k:k+H},\bar{\bm{x}}_{t-1})$ by dropping $\bar{\bm{r}}_{k:k+H}$ in the sequel for notational simplicity. The problem in (13) echoes the MHE problem in (11) in a way as shown below.
@@
$\bar{\bm{w}}_{t}\sim\mathcal{N}\left(\mathbf{0},\bar{\bm{Q}}\right),\quad{\bar{\bm{v}}}_{t}\sim\mathcal{N}\left(\mathbf{0},\bar{\bm{R}}\right),$
@@
Meanwhile, $p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{x}}_{t-1})\sim\mathcal{N}\left(\bar{\bm{f}}(\bar{\bm{x}}_{t-1}),\bar{\bm{Q}}\right)$, which is a degenerate Gaussian distribution as $\bar{\bm{Q}}$ is rank-deficient. By the disintegration theorem [68, Theorem 5.4], $p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{x}}_{t-1})$ collapses to a lower-dimensional distribution based on $p(\bm{w}_{t-1})$, leading to
@@
for $t=k,\ldots,k+H$. Putting together the above, we find that the cost function in (13) is given by
@@ V-A Implicit Importance Sampling
$\mathbb{E}[\bm{g}(\bm{x})\;|\;\bm{y}]\approx\sum_{i=1}^{N}w^{i}\bm{g}(\bm{x}^{i}),\quad w^{i}\propto\frac{1}{N}\frac{p(\bm{x}^{i}\,|\,\bm{y})}{q(\bm{x}^{i})},$ (18)
@@
$w^{i}$ $\propto\left|\frac{d\Gamma(\bm{\xi}^{i})}{d\bm{\xi}^{i}}\right|\cdot\exp[-\min F(\bm{x})+\min\varphi(\bm{\xi})].$
Using the particles along with their weights, we can evaluate $\mathbb{E}[\bm{g}(\bm{x})\;|\;\bm{y}]$ along similar lines in (18).
@@
$\left[\begin{matrix}\bm{x}\cr\bm{y}\end{matrix}\right]\sim\mathcal{N}\left(\left[\begin{matrix}\bm{m}^{\bm{x}}\cr\bm{m}^{\bm{y}}\end{matrix}\right],\left[\begin{matrix}\bm{P}^{\bm{x}}&\bm{P}^{\bm{x}\bm{y}}\cr\bm{\left missing}(\bm{P}^{\bm{x}\bm{y}}\right)^{\top}&\bm{P}^{\bm{y}}\end{matrix}\right]\right),$
then $p(\bm{x}\,|\,\bm{y})\sim\mathcal{N}\left(\tilde{\bm{m}},\tilde{\bm{P}}\right)$, where
@@
$\mathbb{E}[\bm{g}(\bm{x})\;|\;\bm{y}]\approx{1\over N}\sum_{i=1}^{N}\bm{g}(\bm{x}^{i}).$
@@ V-B Implicit Particle Filtering via a Bank of Kalman Filters
Given the system in (12), suppose that we have obtained the particles along with their importance weights $\left\{\bar{\bm{x}}_{k:t-1}^{f,i},w_{k:t-1}^{f,i},\,i=1,\ldots,N\right\}$ at time $t-1$ to form an empirical distribution that approximates $p(\bm{\bar{x}}_{k:t-1}\,|\,\bm{\bar{y}}_{k:t-1})$:
@@
$F^{f}(\bar{\bm{x}}_{t})-\min F^{f}(\bar{\bm{x}}_{t})=\varphi(\bm{\xi}_{t})-\min\varphi(\bm{\xi}_{t}),$ (21)
where $F^{f}(\bar{\bm{x}}_{t})=-\log p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{y}}_{k:t})$ and $\varphi(\bm{\xi}_{t})=-\log p(\bm{\xi}_{t})$.
@@ V-C Implicit Particle Smoothing via a Bank of Kalman Smoothers
$F^{s}(\bar{\bm{x}}_{t})-\min F^{s}(\bar{\bm{x}}_{t})=\varphi(\bm{\xi}_{t})-\min\varphi(\bm{\xi}_{t}),$ (29)
where $F^{s}(\bar{\bm{x}}_{t})=-\log p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{y}}_{k:k+H})$, and $\varphi(\bm{\xi}_{t})$ follows the same definition in (21). As before, an explicit expression for $\bar{\bm{x}}_{t}=\Gamma^{s}(\bm{\xi}_{t})$ will be sought after to solve (29). To achieve this, we impose the following local Gaussian approximations:
@@
2:
3:Set up the virtual system in (9) and (12)
4:for $k=1,2,\ldots$ do
5:     // Forward filtering by Kalman-IPF
6:     Initialize the filtering particles $\bar{\bm{x}}_{k}^{f,i}$ for $i=1,\ldots,N$ with $\bm{\Sigma}_{k}^{f,i}$ and $w_{k}^{f,i}$
7:     for $t=k,k+1,\ldots,k+H$ do
8:         for $i=1,2,\ldots,N$ do
9:              Run KalmanPredict by $\mathcal{UT}$
@@
11:              Draw a highly probable particle $\bm{\xi}_{t}^{i}\sim\mathcal{N}(\bm{0},\bm{I})$, and update the filtering particle
@@
12:         end for
13:         Do resampling if necessary
14:     end for
15:     // Backward smoothing by Kalman-IPS
16:     Initialize the smoothing particles by $\bar{\bm{x}}_{k+H}^{s,i}=\bar{\bm{x}}_{k+H}^{f,i}$ and $\bm{\Sigma}_{k+H}^{s,i}=\bm{\Sigma}_{k+H}^{f,i}$ for $i=1,\ldots,N$
17:     for $t=k+H-1,k+H-2,\ldots,k$ do
18:         for $i=1,2,\ldots,N$ do
19:              Run RTSKalmanUpdate
@@
20:              Draw a highly probable particle $\bm{\xi}_{t}^{i}\sim\mathcal{N}(\bm{0},\bm{I})$, and update the smoothing particle
@@
22:     end for
23:     Compute the final estimate for $\bar{\bm{x}}_{k}$
@@
24:     Export and apply control decisions
25:end for

### v5

@@ I Introduction
In this paper, we develop an alternative framework to perform MPC motion planner design when neural network vehicle models are used. The framework, referred to as model predictive inferential control (MPIC), inherits the idea of optimal motion planning from MPC, but pivots away from using gradient-based optimization as solvers. Instead, it undertakes inference-driven decision-making in planning and attempts to estimate the optimal motion plans while using the driving requirements as the evidence. This profound shift allows us to draw on the substantial work on nonlinear estimation in the literature and utilize various estimation methods to do MPC motion planning. Particularly powerful and attractive among those methods is particle filtering. This technique exploits Monte Carlo sampling to achieve accurate estimation despite strong nonlinearities [7, 8] and thus provides a leverage to treat the difficulty brought by neural networks to motion planning. However, particle filtering in general is computationally expensive, because it requires intensive sampling, especially when the nonlinearity is sophisticated. This motivates us to investigate a new way to do particle filtering in order to execute the MPIC framework. Our approach, by design, will focus on less intensive but more effective sampling to accelerate computation.
@@ II-B Optimal Control via Estimation
The connections between control and estimation have been a fundamental research topic. The seminal work [35] by Rudolf Kalman in 1960 elucidates the duality between the linear quadratic regulator and Kalman smoother for linear systems. Even though such exact duality was long considered hardly generalizable, some illuminating studies manage to formulate nonlinear stochastic optimal control and estimation problems dual to each other [36, 37, 38]. Optimal control computation often requires great amounts of time and memory, but the control-estimation synergy makes it possible to cast control design as explainable estimation problems which are more tractable to solve. The work in [39] proposes stochastic optimal control by variational inference based on Kullback-Leibler divergence minimization, which has initiated a line of inquiry. Especially, a special case of the approach [39] is the path integral control method developed in [40]. This relevance has inspired a body of work known as model predictive path integral control (MPPI) [41, 42]. One can use numerical variational inference methods based on stochastic optimization, Markov Chain Monte Carlo sampling, importance sampling, and others, to compute optimal control for these methods. In [43, 44], stochastic optimal control is handled through recursive Bayesian state estimation based on belief propagation, despite only an approximate relationship between the respective control and estimation problems in this case. However, for deterministic nonlinear systems, optimal control and MPC problems can find equivalent counterparts in recursive Bayesian state estimation problems [45], and particle filtering provides a useful means to deal with these problems due to its power in handling nonlinearity. In our preliminary work [46, 47], we develop constraint-aware particle filtering/smoothing to perform nonlinear MPC with constraints. Closely tied to optimal control, reinforcement learning has found some solutions based on Bayesian estimation [48] and variational inference [49].
@@ II-C MPC of Machine Learning Dynamic Models
The sweeping successes of machine learning has exponentially driven the study of learning-based MPC for dynamic systems [50, 51]. A vast literature has formed around a rich set of topics. We narrow the attention here to the specific subject of MPC of neural networks. Neural networks have a history of being used in the data-driven modeling of dynamic systems. Such models are often called neural state-space (NSS) models, and MPCs for them have garnered many applications [51]. To date, the solvers are mostly based on gradient-based optimization [52, 53, 54, 55, 56, 57]. Gradient-based search, however, will easily lead MPC computation to get stuck in local minima, as NSS models are highly nonlinear and nonconvex. The study in [58] suggests explicitly constructing neural networks that are convex with respect to the input so as to avoid nonconvex optimization, but such input-convex neural networks would have restricted representation capacity. MPC of NSS models is an essential step in model-based reinforcement learning but faces unaffordable computation when applying gradient-based solutions to models based on deep neural networks. In this context, a simple technique is to randomly generate many candidate control decision sequences and then pick the sequence that leads to the minimum cost after being applied to the NSS model [6, 59]. This technique, however, can hardly balance between accuracy and computation. The problem of MPC for NSS models hence still remains widely open to new solutions. On a related note, MPC for systems described by Gaussian processes has recently seen burgeoning studies, e.g.,  [60, 61, 5], and there arise similar challenges.
@@ III-A Neural State-Space Vehicle Modeling
We can construct $\hat{f}$ using different numerical discretization methods. It is often straightforward to use the first-order Euler method, i.e.,
@@ III-B Driving Requirements and Objectives
An autonomous vehicle is expected to operate responsibly within traffic, demonstrating safety, ethics, and predictability in its maneuvers. To this end, it should adhere to some driving requirements and constraints. Specifically, the vehicle (referred to as the ego vehicle or EV) must avoid collision with other traffic participants (referred to as obstacle vehicles or OVs), remain within the boundaries of the road, operate in its practical actuation limits, and follow some nominal guidance (in path, velocity, etc.) generated by a higher-level decision maker.
@@
$\mathcal{S}_{k}=\mathcal{S}\left(\bm{x}_{k},\mathcal{B}\right),\ \ \mathcal{O}_{k}^{i}=\mathcal{O}\left(\bm{x}_{k}^{\mathrm{OV},i},\mathcal{B}\right),\ i=1,2,\ldots,N_{O}.$
@@
Vehicle Actuation Limits: The EV’s actuation $\bm{u}_{k}$, including the acceleration and steering angle, is subject to physical limits during the maneuver. We thus constrain $\bm{u}_{k}$ in motion planning as follows:
@@ III-C Motion Planning Synthesis
Having laid out the driving requirements and objectives, we are now ready to synthesize the motion planning problem. To begin with, we express the constraints in (2)-(5) notationally as
@@
$\min\quad$ $J\left(\bm{x}_{k:k+H},\bm{u}_{k:k+H},\Delta\bm{u}_{k:k+H}\right),$ (8a)
@@ IV-A MPC Through the Lens of State Estimation
$\left\{\begin{aligned} \bm{x}_{t+1}&=\hat{f}(\bm{x}_{t},\bm{u}_{t}),\\ \bm{u}_{t+1}&=\bm{u}_{t}+\Delta\bm{u}_{t+1},\\ \Delta\bm{u}_{t+1}&=\bm{w}_{t},\\ \bm{y}_{\bm{x},t}&=\bm{x}_{t}+\bm{v}_{\bm{x},t},\\ \bm{y}_{\bm{u},t}&=\bm{u}_{t}+\bm{v}_{\bm{u},t},\\ y_{g,t}&=\sum_{j=1}^{m}\psi\left(g_{j}\left(\bm{x}_{t},\bm{u}_{t},\Delta\bm{u}_{t}\right)\right)+\varepsilon_{t},\end{aligned}\right$. (9)
@@ IV-B MPIC Based on Bayesian State Estimation
$\bar{\bm{f}}\left(\bar{\bm{x}}_{t}\right)=\begin{bmatrix}\hat{f}(\bm{x}_{t},\bm{u}_{t})\\ \bm{u}_{t}\\ \bm{0}\end{bmatrix},\;\bar{\bm{h}}(\bar{\bm{x}}_{t})=\left[\begin{matrix}\bm{x}_{t}\cr\bm{u}_{t}\cr\displaystyle\sum_{j=1}^{m}\psi\left(g_{j}(\bm{x}_{t},\bm{u}_{t},\Delta\bm{u}_{t})\right)\end{matrix}\right].$
@@
We will use $p(\bar{\bm{x}}_{k:k+H}\,|\,\bar{\bm{y}}_{k:k+H},\bar{\bm{x}}_{k-1})$ by dropping $\bar{\bm{r}}_{k:k+H}$ in the sequel for notational simplicity. The problem in (13) echoes the MHE problem in (11) in a way as shown below.
@@
$\bar{\bm{w}}_{t}\sim\mathcal{N}\left(\mathbf{0},\bar{\bm{Q}}\right),\quad{\bar{\bm{v}}}_{t}\sim\mathcal{N}(\mathbf{0},\bar{\bm{R}}),$
@@
Meanwhile, $p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{x}}_{t-1})\sim\mathcal{N}\left(\bar{\bm{f}}(\bar{\bm{x}}_{t-1}),\bar{\bm{Q}}\right)$, which is a degenerate Gaussian distribution as $\bar{\bm{Q}}$ is rank-deficient. By the disintegration theorem [67, Theorem 5.4], $p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{x}}_{t-1})$ collapses to a lower-dimensional distribution based on $p(\bm{w}_{t-1})$, leading to
@@
for $t=k,\ldots,k+H$.
Putting together the above, we find that the cost function in (13) is given by
@@ V-A Implicit Importance Sampling
$\mathbb{E}\left[\bm{g}(\bm{x})\;|\;\bm{y}\right]\approx\sum_{i=1}^{N}w^{i}\bm{g}(\bm{x}^{i}),\quad w^{i}\propto\frac{1}{N}\frac{p(\bm{x}^{i}\,|\,\bm{y})}{q(\bm{x}^{i})},$ (18)
@@
$w^{i}$ $\propto\left|\frac{d\Gamma(\bm{\xi}^{i})}{d\bm{\xi}^{i}}\right|\cdot\exp\left[-\min F(\bm{x})+\min\varphi(\bm{\xi})\right].$
Using the particles along with their weights, we can evaluate $\mathbb{E}\left[\bm{g}(\bm{x})\;|\;\bm{y}\right]$ along similar lines in (18).
@@
$\left[\begin{matrix}\bm{x}\cr\bm{y}\end{matrix}\right]\sim\mathcal{N}\left(\left[\begin{matrix}\bm{m}^{\bm{x}}\cr\bm{m}^{\bm{y}}\end{matrix}\right],\left[\begin{matrix}\bm{P}^{\bm{x}}&\bm{P}^{\bm{x}\bm{y}}\cr\bm{\left.}(\bm{P}^{\bm{x}\bm{y}}\right)^{\top}&\bm{P}^{\bm{y}}\end{matrix}\right]\right),$
[…]
@@
$\mathbb{E}\left[\bm{g}(\bm{x})\;|\;\bm{y}\right]\approx{1\over N}\sum_{i=1}^{N}\bm{g}(\bm{x}^{i}).$
@@ V-B Implicit Particle Filtering via a Bank of Kalman Filters
Given the system in (12), suppose that we have obtained the particles along with their importance weights $\left\{\bar{\bm{x}}_{k:t-1}^{f,i},w_{k:t-1}^{f,i},\,i=1,\ldots,N\right\}$ at time t1t-1 to form an empirical distribution that approximates $p(\bm{\bar{x}}_{k:t-1}\,|\,\bm{\bar{y}}_{k:t-1})$:
@@
$F^{f}\left(\bar{\bm{x}}_{t}\right)-\min F^{f}\left(\bar{\bm{x}}_{t}\right)=\varphi(\bm{\xi}_{t})-\min\varphi(\bm{\xi}_{t}),$ (21)
where $F^{f}\left(\bar{\bm{x}}_{t}\right)=-\log p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{y}}_{k:t})$ and $\varphi(\bm{\xi}_{t})=-\log p(\bm{\xi}_{t})$.
@@ V-C Implicit Particle Smoothing via a Bank of Kalman Smoothers
$F^{s}\left(\bar{\bm{x}}_{t}\right)-\min F^{s}\left(\bar{\bm{x}}_{t}\right)=\varphi(\bm{\xi}_{t})-\min\varphi(\bm{\xi}_{t}),$ (29)
where $F^{s}\left(\bar{\bm{x}}_{t}\right)=-\log p(\bar{\bm{x}}_{t}\,|\,\bar{\bm{y}}_{k:k+H})$, and $\varphi(\bm{\xi}_{t})$ follows the same definition in (21). As before, an explicit expression for $\bar{\bm{x}}_{t}=\Gamma^{s}(\bm{\xi}_{t})$ will be sought after to solve (29). To achieve this, we impose the following local Gaussian approximations:
@@
2:Set up the virtual system in (9)
3:for $k=1,2,\ldots$ do
4:  // Forward filtering by Kalman-IPF
5:  Initialize the filtering particles $\bar{\bm{x}}_{k}^{f,i}$ for $i=1,\ldots,N$ with $\bm{\Sigma}_{k}^{f,i}$ and $w_{k}^{f,i}$
6:  for $t=k,k+1,\ldots,k+H$ do
7:   for $i=1,2,\ldots,N$ do
8:     Run KalmanPredict by $\mathcal{UT}$
@@
10:     Draw a highly probable particle $\bm{\xi}_{t}^{i}\sim\mathcal{N}(\bm{0},\bm{I})$, and update the filtering particle
@@
11:   end for
12:   Do resampling if necessary
13:  end for
14:  // Backward smoothing by Kalman-IPS
15:  Initialize the smoothing particles by $\bar{\bm{x}}_{k+H}^{s,i}=\bar{\bm{x}}_{k+H}^{f,i}$ and $\bm{\Sigma}_{k+H}^{s,i}=\bm{\Sigma}_{k+H}^{f,i}$ for $i=1,\ldots,N$
16:  for $t=k+H-1,k+H-2,\ldots,k$ do
17:   for $i=1,2,\ldots,N$ do
18:     Run RTSKalmanUpdate
@@
19:     Draw a highly probable particle $\bm{\xi}_{t}^{i}\sim\mathcal{N}(\bm{0},\bm{I})$, and update the smoothing particle
@@
22:  Compute the final estimate for $\bar{\bm{x}}_{k}$
@@
23:  Export and apply control decisions
24:end for